# Style Amnesia: Investigating Speaking Style Degradation and Mitigation in Multi-Turn Spoken Language Models

### v1

@@ Abstract
In this paper, we show that when spoken language models (SLMs) are instructed to speak in a specific speaking style at the beginning of a multi-turn conversation, they cannot maintain the required speaking styles after several turns of interaction; we refer to this as the style amnesia of SLMs. We focus on paralinguistic speaking styles, including emotion, accent, volume, and speaking speed. We evaluate three proprietary and two open-source SLMs, demonstrating that none of these models can maintain a consistent speaking style when instructed to do so. We further show that when SLMs are asked to recall the style instruction in later turns, they can recall the style instruction, but they fail to express it throughout the conversation. We also show that explicitly asking the model to recall the style instruction can partially mitigate style amnesia. In addition, we examine various prompting strategies and find that SLMs struggle to follow the required style when the instruction is placed in system messages rather than user messages, which contradicts the intended function of system prompts.
@@ 1 Introduction
Spoken language models (SLMs) can take speech input and generate speech responses. Unlike text-only large language models (LLMs), SLMs integrate audio encoders and vocoders Kong et al. (2020) to support end-to-end (E2E) speech understanding and generation (Défossez et al., 2024; Zeng et al., 2024; Fang et al., 2025; Wu et al., 2025; Huang et al., 2025). While this E2E design introduces additional challenges beyond text processing, such as handling paralinguistic features and non-textual information, current SLMs have demonstrated the capability to detect user attributes like emotion, accent, and gender. By leveraging these acoustic cues, SLMs can adjust their outputs to produce more contextually appropriate responses (Wu et al., 2025; Google, 2025b). Beyond passive perception, SLMs are also capable of actively following the user-specified speaking style in single-turn interactions (Zeng et al., 2024; Wu et al., 2025; Chiang et al., 2025; Zhan et al., 2025).
@@
  • We show that explicitly prompting SLMs to recall the initial style instruction can alleviate style degradation, but it does not entirely resolve the issue.
@@ 2.1 Context Loss in Multi-Turn Dialogue
Although LLMs support multi-turn interactions, recent studies have shown that their performance often degrades as dialogues become longer Kwan et al. (2024); Han et al. (2025); Qamar et al. (2025); Li et al. (2025); Laban et al. (2025). These studies indicate the shortcomings of current LLMs, such as difficulties with multi-turn instruction following Han et al. (2025); Li et al. (2025), loss of information in long dialogue Kwan et al. (2024); Qamar et al. (2025), and errors caused by fragmented information across turns Laban et al. (2025).
In terms of SLMs evaluation, SpokenWOZ Si et al. (2023) highlights the difficulty of aggregating fragmented information across turns. $C^{3}$ Ma et al. (2025) focuses on bilingual spoken dialogue and evaluates models’ capabilities in handling complex multi-turn conversational phenomena, including omission and coreference. ContextDialog Kim et al. (2025), on the other hand, demonstrates that SLMs exhibit significantly lower memorization performance in long dialogue compared to their text-based counterparts. Overall, these works assess the ability of SLMs to retain and utilize information in long spoken dialogues but leave the expressiveness of generated speech unevaluated.
@@
SpokenWOZ Si et al. (2023)
C3 Ma et al. (2025)
ContextDialog Kim et al. (2025)
Vstyle Zhan et al. (2025)
Game-Time Chang et al. (2025)
StyleSet Chiang et al. (2025)
URO-Bench Yan et al. (2025)
VocalBench Liu et al. (2025)
VoxDialogue Cheng et al. (2025)
Multi-Bench Deng et al. (2025)
@@ 2.2 Speaking Style Assessment
Beyond basic intelligibility and audio quality evaluation Saeki et al. (2022); Fang et al. (2025), a growing body of work has started to focus on the expressiveness of SLMs Zeng et al. (2024); Wu et al. (2025). Vstyle Zhan et al. (2025) evaluates whether SLMs can adapt their voice style according to the given instruction. Game-Time Chang et al. (2025) focuses on the temporal control of SLMs. Besides, Chiang et al. (2025) demonstrate that Large Audio-Language Models (LALMs) are capable of serving as an automatic judge to evaluate the voice style. Their experiment shows high agreement between human evaluation results and those from Gemini-2.5 Pro Comanici et al. (2025).
Although previous works have evaluated the expressiveness of SLMs, voice style consistency in multi-turn interactions remains largely underexplored. URO-bench Yan et al. (2025), VocalBench Liu et al. (2025), and VoxDialogue Cheng et al. (2025) use predefined dialogues as the first k1k-1 turns and ask SLMs to generate the $k$-th response, which does not support turn-level analysis. In contrast to approaches based on predefined multi-turn dialogues, we adopt a model-based user simulator to interact with the evaluated SLMs. We argue that this setting more closely reflects real-world conversational scenarios.
A concurrent work, Multi-Bench Deng et al. (2025), also evaluates SLMs in multi-turn interactive dialogues. However, their evaluation aggregates performance into a single global score to assess overall emotional intelligence. In contrast, our turn-level analysis provides fine-grained and informative insights into how inconsistencies emerge across a wide range of speaking styles, including emotion, accent, speed, and volume. A comparison with prior work is summarized in Table 1.
@@ 3.2.2 Dialogue Topics
In the first user turn of the dialogue, we provide a conversation opener for each dialogue. The conversation opener is used to control the topic that the SLM should talk about during each evaluation, ensuring that different SLMs are evaluated under a common conversational setting. We select the topic from Soda Kim et al. (2023), a large-scale dataset of social interaction dialogues. The details about how we select the topics from Soda are shown in Appendix B. Eventually, we collect 100 diverse conversation openers to initiate the conversations. We manually inspect the selected topics and filter out improper topics, such as queries about personal preferences that are unsuitable for machine conversations. An example conversation opener is shown in Figure 2 “How can I fight off sleepiness?”
@@ 3.2.3 Multi-Turn Interaction Using a User Simulator
At the core of our evaluation framework is the multi-turn interaction between the evaluated SLM and a user. To enable back-and-forth interactions with an evaluated SLM, we build a user simulator with a cascade SLM. The cascade SLM is a speech-in-speech-out system composed of an ASR, a text-only LLM, and a TTS model. The ASR model takes the speech input from the evaluated SLM and transcribes it into text. [^1] Most existing SLMs output both speech and its corresponding transcription at the same time. In this case, we directly use the text output from the evaluated SLM and skip the ASR module in the user simulator. A text-only LLM, powered by GPT-5 mini OpenAI (2025), takes the text input and generates a text response based on the transcription produced by the evaluated SLM. Last, the TTS model, GPT-4o mini TTS OpenAI (2024), converts the text output from the LLM into speech and sends it back to the evaluated SLM to continue the conversation.
@@ 3.3 Evaluation Metrics
(1) First Turn IF Rate $IF_{1}$.  This metric serves as a reference to evaluate the capability of the SLM to follow the specific style in the first turn. If an SLM fails to achieve a reasonable $IF_{1}$, we may not expect it to perform better in later turns.
@@
Aside from measuring the speaking style, we also measure the semantic coherence of the dialogue to ensure that the content is valid using LLM-as-a-judge (Chiang and Lee, 2023). Since the semantic coherence of the dialogue is not our primary focus, we leave the details in Appendix D.3. Overall, all SLMs we evaluate maintain reasonable semantic consistency across turns.
@@ 3.3.1 Automatic Judges for Styles
Emotion We use Emotion2vec-Large Ma et al. (2024) to predict the emotion of $o_{i,j}$. Although it is a 9-class model, we focus on the probabilities for happiness, sadness, anger, and neutral. We take the maximum probability among these four categories as the final prediction and verify whether it aligns with $s$. If the prediction is aligned, the indicator function $\mathbbm{1}(\cdot)$ is 1; otherwise, it is 0.
Accent We use Voxlect-English-Dialect-Whisper-Large-v3 Feng et al. (2025a) to predict the accent of $o_{i,j}$. From this 16-class model, we extract the probabilities for North American and Indian English, taking the maximum as the predicted accent of $o_{i,j}$. If it is aligned with $s$, $\mathbbm{1}(\cdot)$ is 1; otherwise $\mathbbm{1}(\cdot)$ is 0.
Volume We measure volume using Loudness Units Full Scale (LUFS) with PyLoudnorm Steinmetz and Reiss (2021). Since subjective terms like “loud” or “quiet” lack universal definitions, we adopt a relative evaluation standard. We first establish a baseline by querying the SLM with a neutral instruction: “You are a text-to-speech model. Please read the given text at a normal volume without adding or omitting anything,” and generate another speech $o^{\prime}_{i,j}$. $\mathbbm{1}(\cdot)$ is 1 if the LUFS of $o_{i,j}$ is higher (for loud instruction) or lower (for quiet instruction) than $o^{\prime}_{i,j}$; otherwise $\mathbbm{1}(\cdot)$ is 0.
Speed We measure speaking rate using Words Per Minute (WPM) with Parakeet TDT v2 Nvidia (2025). Similar to volume, we compare the SLM generated speech $o_{i,j}$ with a TTS-generated speech $o^{\prime}_{i,j}$ using the same model to determine whether $o_{i,j}$ is faster (or slower) than $o^{\prime}_{i,j}$.
Model Emotion Accent
Happiness Neutral Anger Sadness North American Indian
$IF_{1}$ $\uparrow$ $D$ $\downarrow$ $IF_{1}$ $\uparrow$ $D$ $\downarrow$ $IF_{1}$ $\uparrow$ $D$ $\downarrow$ $IF_{1}$ $\uparrow$ $D$ $\downarrow$ $IF_{1}$ $\uparrow$ $D$ $\downarrow$ $IF_{1}$ $\uparrow$ $D$ $\downarrow$
Cascaded Baseline 85.0 3.0 73.0 0.0 17.0 2.3 62.0 2.0 100.0 0.3 100.0 0.3
Gemini Live 91.0 1.3 64.0 0.0 24.0 14.0 72.0 21.3 100.0 0.0 100.0 0.0
GPT-4o 82.1 0.0 58.8 0.7 30.5 13.7 78.0 26.7 100.0 0.0 100.0 1.7
GPT-4o mini 89.0 2.3 45.0 11.0 39.0 34.7 85.0 65.3 100.0 0.0 89.0 49.7
Qwen2.5-Omni 71.0 0.0 41.0 22.7 5.0 3.7 17.0 14.3 100.0 0.3 0.0 0.0
Step-Audio 2 mini 100.0 2.0 7.0 3.3 1.0 1.0 17.0 14.0 66.0 0.0 32.0 5.0
Model Volume Speed
Loud Quiet Fast Slow
$IF_{1}\uparrow$ $D\downarrow$ $IF_{1}\uparrow$ $D\downarrow$ $IF_{1}\uparrow$ $D\downarrow$ $IF_{1}\uparrow$ $D\downarrow$
Cascaded Baseline 96.0 0.0 99.0 1.3 100.0 0.7 100.0 0.7
Gemini Live 57.0 0.7 95.0 7.3 99.0 9.7 99.0 0.7
GPT-4o 67.0 6.3 92.7 0.0 89.0 3.3 100.0 8.1
GPT-4o mini 77.0 9.3 100.0 1.0 87.0 4.0 99.0 20.9
Qwen2.5-Omni 38.0 1.3 69.0 7.0 47.0 14.3 40.0  0.0
Step-Audio 2 mini 46.0 0.7 56.0 3.0 89.0 27.0 81.0 27.0
Table 2: The First Turn IF Rate $IF_{1}$ and Degradation Rate $D$ across different speaking styles.
@@ 4.1 Experimental Setup
We evaluate three proprietary SLMs: GPT-4o 222gpt-4o-audio-preview-2025-06-03OpenAI (2024), GPT-4o mini 333gpt-4o mini-audio-preview-2024-12-17OpenAI (2024), Gemini Live 444gemini-2.5-flash-native-audio-preview-09-2025Google (2025b), and two open-source end-to-end SLMs: Qwen2.5-Omni Xu et al. (2025) and Step-Audio 2 mini Wu et al. (2025). We select these two open-source SLMs because they are representative models and they provide official vLLM implementations Kwon et al. (2023) to speed up the inference process.
Apart from these E2E SLMs, we introduce a cascaded SLM baseline comprised of GPT-5 mini and Gemini-TTS Google (2025a). The TTS in the baseline receives the style instruction in each turn. The performance of the cascaded baseline serves as the performance upper-bound. The hyperparameter choices of each model are provided in Appendix C.1.
@@ 4.2 Results
We report the First Turn IF Rate $IF_{1}$ and the Degradation Rate $D$ in Table 2.555GPT-4o and GPT-4o mini sometimes return only text transcription without speech. When this occurs, we query the models up to three times, but some samples still fail. We discuss this issue in Appendix D.6. In the following section, all metrics are computed only on successful cases. We also provide the visualizations of style amnesia in Figure 3, and IF Rate for each turn in Appendix D.1.
For the Emotion speaking style, the Cascaded Baseline demonstrates consistent IF rate across turns, with degradation within $\pm 3.0\%$. In contrast, Gemini Live and GPT-4o show a degradation rate of approximately 10% for Anger and Sadness. GPT-4o mini exhibits even larger degradation, reaching 34.7% and 65.3% for Anger and Sadness, respectively. Qwen2.5-Omni and Step-Audio 2 mini also show around 14.0% degradation for Sadness. However, their degradation rates for Anger are only -3.7% and -1.0%, because both models already struggle to perform this style effectively in the first turn.
@@
Figure 4: The difference of First Turn IF Rate when instructions are placed in system and user messages. First Turn IFRate refers to the IF rate in first turn.
We also observe style degradation when SLMs are instructed to maintain Volume and Speed. Regarding Volume, results indicate that speaking loudly is generally more challenging for SLMs than speaking quietly, even in their first turn. Despite this, models capable of volume adjustment still exhibit some degradation. In terms of Speed, most SLMs demonstrate reasonable control in the first turn but show significant degradation over time. An exception is Qwen2.5-Omni, which fails to control speed even in the first turn, as evidenced by an $IF_{1}$ score below 50%, a value comparable to a random baseline in our pairwise comparison setting.
Based on these experiments, we demonstrate that SLMs are prone to losing control of speaking style in multi-turn conversations. We further conduct additional experiments on Step-Audio 2 mini in Appendix D.4, which show that prompt paraphrasing does not alleviate the style degradation issue.

### v2

@@ Abstract
In this paper, we show that when spoken language models (SLMs) are instructed to speak in a specific speaking style at the beginning of a multi-turn conversation, they cannot maintain the required speaking styles after several turns of interaction; we refer to this as the style amnesia of SLMs. We focus on paralinguistic speaking styles, including emotion, accent, volume, and speaking speed. We evaluate three proprietary and two open-source SLMs, demonstrating that none of these models can maintain a consistent speaking style when instructed to do so. We further show that when SLMs are asked to recall the style instruction in later turns, they can recall the style instruction, but they fail to express it throughout the conversation. We also show that explicitly asking the model to recall the style instruction can partially mitigate style amnesia. In addition, we examine various prompting strategies and find that SLMs struggle to follow the required style when the instruction is placed in system messages rather than user messages, which contradicts the intended function of system messages.
@@ 1 Introduction
Spoken language models (SLMs) can take speech input and generate speech responses. Unlike text-only large language models (LLMs), SLMs integrate audio encoders and vocoders kong2020hifi to support end-to-end (E2E) speech understanding and generation (defossez2024moshi; zeng2024glm; fangllama; wu2025step; huang2025step). While this E2E design introduces additional challenges beyond text processing, such as handling paralinguistic features and non-textual information, current SLMs have demonstrated the capability to detect user attributes like emotion, accent, and gender. By leveraging these acoustic cues, SLMs can adjust their outputs to produce more contextually appropriate responses (wu2025step; GeminiLiveUpdate2025). Beyond passive perception, SLMs are also capable of actively following the user-specified speaking style in single-turn interactions (zeng2024glm; wu2025step; chiang2025audio; zhan2025vstyle).
@@
  • We show that explicitly prompting SLMs to recall the initial style instruction can alleviate style amnesia, but it does not entirely resolve the issue.
@@ 2.1 Context Loss in Multi-Turn Dialogue
Although LLMs support multi-turn interactions, recent studies have shown that their performance often degrades as dialogues become longer kwan-etal-2024-mt; han-etal-2025-language; qamar-etal-2025-llms; li-etal-2025-structflowbench; laban2025llms. These studies indicate the shortcomings of current LLMs, such as difficulties with multi-turn instruction following han-etal-2025-language; li-etal-2025-structflowbench, loss of information in long dialogue kwan-etal-2024-mt; qamar-etal-2025-llms, and errors caused by fragmented information across turns laban2025llms.
In terms of SLMs evaluation, SpokenWOZ si2023spokenwoz highlights the difficulty of aggregating fragmented information across turns. $C^{3}$ ma2025c3 focuses on bilingual spoken dialogue and evaluates models’ capabilities in handling complex multi-turn conversational phenomena, including omission and coreference. ContextDialog kim-etal-2025-voice-assistant, on the other hand, demonstrates that SLMs exhibit significantly lower memorization performance in long dialogue compared to their text-based counterparts. Overall, these works assess the ability of SLMs to retain and utilize information in long spoken dialogues but leave the expressiveness of generated speech unevaluated.
@@
SpokenWOZ si2023spokenwoz
C3 ma2025c3
ContextDialog kim-etal-2025-voice-assistant
Vstyle zhan2025vstyle
Game-Time chang2025game
StyleSet chiang2025audio
URO-Bench yan-etal-2025-uro
VocalBench liu2025vocalbench
VoxDialogue cheng2025voxdialogue
Multi-Bench deng2025multi
@@ 2.2 Speaking Style Assessment
Beyond basic intelligibility and audio quality evaluation saeki2022utmos; fangllama, a growing body of work has started to focus on the expressiveness of SLMs zeng2024glm; wu2025step. For example, Vstyle zhan2025vstyle evaluates whether SLMs can adapt their voice style according to the given instruction. Game-Time chang2025game focuses on the temporal control of SLMs. Besides, chiang2025audio demonstrate that Large Audio-Language Models (LALMs) are capable of serving as an automatic judge to evaluate the voice style. Their experiment shows high agreement between human evaluation results and those from Gemini-2.5 Pro comanici2025gemini.
Although previous works have evaluated the expressiveness of SLMs, voice style consistency in multi-turn interactions remains largely underexplored. URO-bench yan-etal-2025-uro, VocalBench liu2025vocalbench, and VoxDialogue cheng2025voxdialogue use predefined dialogues as the first k1k-1 turns and ask SLMs to generate the $k$-th response, which does not support turn-level analysis. In contrast to approaches based on predefined multi-turn dialogues, we adopt a model-based user simulator to interact with the evaluated SLMs. We argue that this setting more closely reflects real-world conversational scenarios.
A concurrent work, Multi-Bench deng2025multi, also evaluates SLMs in multi-turn interactive dialogues. However, their evaluation aggregates performance into a single global score to assess overall emotional intelligence. In contrast, our turn-level analysis provides fine-grained and informative insights into how inconsistencies emerge across a wide range of speaking styles, including emotion, accent, speed, and volume. A comparison with prior work is summarized in Table 1.
@@ 3.2.2 Dialogue Topics
In the first user turn of the dialogue, we provide a conversation opener for each dialogue. The conversation opener is used to control the topic that the SLM should talk about during each evaluation, ensuring that different SLMs are evaluated under a common conversational setting. We select the topic from Soda kim2023soda, a large-scale dataset of social interaction dialogues. The details about how we select the topics from Soda are shown in Appendix B. Eventually, we collect 100 diverse conversation openers to initiate the conversations. We manually inspect the selected topics and filter out improper topics, such as queries about personal preferences that are unsuitable for machine conversations. An example conversation opener is shown in Figure 2 “How can I fight off sleepiness?”
@@ 3.2.3 Multi-Turn Interaction Using a User Simulator
At the core of our evaluation framework is the multi-turn interaction between the evaluated SLM and a user. To enable back-and-forth interactions with an evaluated SLM, we build a user simulator with a cascade SLM. The cascade SLM is a speech-in-speech-out system composed of an ASR, a text-only LLM, and a TTS model. The ASR model takes the speech input from the evaluated SLM and transcribes it into text. [^1] Most existing SLMs output both speech and its corresponding transcription at the same time. In this case, we directly use the text output from the evaluated SLM and skip the ASR module in the user simulator. A text-only LLM, powered by GPT-5 mini OpenAIGPT5mini2024, takes the text input and generates a text response based on the transcription produced by the evaluated SLM. Last, the TTS model, GPT-4o mini TTS OpenAIGPT4oMini2024, converts the text output from the LLM into speech and sends it back to the evaluated SLM to continue the conversation.
@@ 3.3 Evaluation Metrics
(1) First-turn IF rate $IF_{1}$.  This metric serves as a reference to evaluate the capability of the SLM to follow the specific style in the first turn. If an SLM fails to achieve a reasonable $IF_{1}$, we may not expect it to perform better in later turns.
@@
Aside from measuring the speaking style, we also measure the semantic coherence of the dialogue to ensure that the content is valid using LLM-as-a-judge (chiang-lee-2023-large). Since the semantic coherence of the dialogue is not our primary focus, we leave the details in Appendix D.3. Overall, all SLMs we evaluate maintain reasonable semantic consistency across turns.
@@ 3.3.1 Automatic Judges for Styles
Emotion We use Emotion2vec-Large ma2024emotion2vec to predict the emotion of $o_{i,j}$. Although it is a 9-class model, we focus on the probabilities for happiness, sadness, anger, and neutral. We take the maximum probability among these four categories as the final prediction and verify whether it aligns with $s$. If the prediction is aligned, the indicator function $\mathbbm{1}(\cdot)$ is 1; otherwise, it is 0.
Accent We use Voxlect-English-Dialect-Whisper-Large-v3 feng2025voxlect to predict the accent of $o_{i,j}$. From this 16-class model, we extract the probabilities for North American and Indian English, taking the maximum as the predicted accent of $o_{i,j}$. If it is aligned with $s$, $\mathbbm{1}(\cdot)$ is 1; otherwise $\mathbbm{1}(\cdot)$ is 0.
Volume We measure volume using Loudness Units Full Scale (LUFS) with PyLoudnorm steinmetz2021pyloudnorm. Since subjective terms like “loud” or “quiet” lack universal definitions, we adopt a relative evaluation standard. We first establish a baseline by querying the SLM with a neutral instruction: “You are a text-to-speech model. Please read the given text at a normal volume without adding or omitting anything,” and generate another speech $o^{\prime}_{i,j}$. $\mathbbm{1}(\cdot)$ is 1 if the LUFS of $o_{i,j}$ is higher (for loud instruction) or lower (for quiet instruction) than $o^{\prime}_{i,j}$; otherwise $\mathbbm{1}(\cdot)$ is 0.
Speed We measure speaking rate using Words Per Minute (WPM) with Parakeet TDT v2 nvidiaasr2025. Similar to volume, we compare the SLM generated speech $o_{i,j}$ with the TTS-generated speech $o^{\prime}_{i,j}$ using the same model to determine whether $o_{i,j}$ is faster (or slower) than $o^{\prime}_{i,j}$.
Figure 3: The first-turn IF rate $IF_{1}$ and degradation rate $D$ across different speaking styles.
@@ 4.1 Experimental Setup
We evaluate three proprietary SLMs: GPT-4o222gpt-4o-audio-preview-2025-06-03openai2024gpt4ocard, GPT-4o mini333gpt-4o mini-audio-preview-2024-12-17OpenAIGPT4oMini2024, Gemini Live444gemini-2.5-flash-native-audio-preview-09-2025GeminiLiveUpdate2025, and two open-source end-to-end SLMs: Qwen2.5-Omni xu2025qwen2 and Step-Audio 2 mini wu2025step. We select these two representative open-source SLMs because they provide official vLLM implementations kwon2023efficient to speed up the inference process.
Apart from these E2E SLMs, we introduce a cascaded SLM baseline comprised of GPT-5 mini and Gemini-TTS geminitts2025. The TTS in the baseline receives the style instruction in each turn. The performance of the cascaded baseline serves as the performance upper-bound. The hyperparameter choices of each model are provided in Appendix C.1.
@@ 4.2 Results
We report the first-turn IF rate $IF_{1}$ and the regradation rate $D$ in Figure 3.555GPT-4o and GPT-4o mini sometimes return only text transcription without speech. When this occurs, we query the models up to three times, but a few samples still fail. We discuss this issue in Appendix D.5. In the following section, all metrics are computed only on successful cases. We also provide the illustration of style amnesia in Figures 1 and 4, and IF rate for each turn in Appendix D.1.
For the Emotion speaking style, the Cascaded Baseline demonstrates consistent IF rate across turns, with degradation within $3.0\%$. In contrast, Gemini Live and GPT-4o show a degradation rate ranging from 13.7% to 26.7% for Anger and Sadness. GPT-4o mini exhibits even larger degradation, reaching 34.7% and 65.3% for Anger and Sadness, respectively. Qwen2.5-Omni and Step-Audio 2 mini also show around 14.0% degradation for Sadness. However, their degradation rates for Anger are only 3.7% and 1.0%, because both models already struggle to perform this style effectively in the first turn.
@@
Figure 5: The difference of first-turn IF $IF_{1}$ rate when instructions are placed in system and user messages.
We also observe style amnesia when SLMs are instructed to maintain Volume and Speed. Regarding Volume, results indicate that speaking loudly is generally more challenging for SLMs than speaking quietly, even in their first turn. Despite this, models capable of volume adjustment still exhibit some degradation. In terms of Speed, most SLMs demonstrate reasonable control in the first turn but show significant degradation over time. An exception is Qwen2.5-Omni, which fails to control speed even in the first turn, as evidenced by an $IF_{1}$ score below 50%, a value comparable to a random baseline in our pairwise comparison setting.
Based on these experiments, we demonstrate that SLMs are prone to losing control of speaking style in multi-turn conversations.